D . The avera ge price per gallon of regular gasoline was reported to be $22.34 in Ohio . Use this price as the population m ean,, and assume that the popula tion standard deviation is $00.16 .

7 . For a random sample of 64 service stations,, find the standard deviation of the sampling distribution of the sample mean (uuse exactly two de c imal places)) .

8 . What is the probability that the mean price for a sample of 64 s ervice stations is between $22.31 and $ 2.36 ? (PP rovide your answer using exactly four decimal places))   
Answer #1

Given,

\mu = 22.34 , \sigma = 0.16

7)

Standard deviation of sampling distribution of sample mean is \sigma\bar{x} .

= \sigma / sqrt(n)

= 0.16 / sqrt(64)

= 0.02

8)

The central limit theorem is

P( \bar{x} < x) = P( Z < \bar{x} - \mu /  \sigma / sqrt(n) )

P( 22.31 < \bar{x} < 22.36) = P( \bar{x} < 22.36) - P( \bar{x} < 22.31)

= P( Z < 22.36 - 22.34 / 0.02) - P( Z < 22.31 - 22.34 / 0.02)

= P( Z < 1) - P( Z < -1.5)

= P( Z < 1 ) - ( 1 - P( Z < 1.5) )

= 0.8413 - ( 1 - 0.9332)

= 0.7745